Journal of nothing First and Last Volume Copyright 1998 by Myself

Some serial number

Sample Title

Myself Author

My Department Address ZIP

Second Author and Some Friend

The other place Elsewhere

Text of the abstract

Here goes the text of the article.

You can have some text appear differently in each mode, or only appear in some mode(s). You can also modify the appearance of tables, figures, or equations between modes of compilation. Here’s an unexciting example:

This is not man mode. You can also use the three-way switch: the current mode is jou.

Sectioning commands are used as in regular L^{A}TEX. It is very important that you use these sectioning commands. Here’s a typical example for an experimental psychology paper, following the (untitled) introductory section:

Experiment 1

Introduction to this experiment. When referring to it, type Experiment 1 (with capital E and a tie to the number so they won’t separate). However, when mentioning the first experiment the “e” is lowercase.

Method

You may need some general remarks on method here. Then use the standard parts.

Subjects. Description of subjects.

Apparatus. Sometimes it is necessary to give the apparatus a special section.

Stimuli. What the subjects saw, heard, or felt.

Procedure. What happened to the poor subjects.

Results

Now the numbers: What your measurements showed and all the statistics of it.

Discussion

If your results mean anything, this is the place to talk about it. In case there is little to discuss before the final section (e.g.,

Thanks to people and grants, and your new address maybe.

General Discussion

at the end) you may combine results and discussion in a single

Results and Discussion

before you go on to the next

Experiment 2

and so on.

In case you have more than these three levels of headings (sectioning), you can use the command e.g., for four levels). See the APA manual for how to handle more or fewer heading levels (the appearance is, naturally, taken care of by apa.cls and you only need concern yourself with the conceptual divisions). In case of more than three levels you’ll need the

(. for the fourth level) and

( for the fifth level) sectioning commands of L^{A}TEX.

Floating things Figures

A figure can be included in any way you like, but this class provides a way to deal automatically with fitting encapsulated postscript (EPS) files without worrying about dimensions or mode. Start your figure or figure* environment as usual (see the LaTeX book) and then instead of “epsfig or “includegraphics use the new command “fitfigure:

An example of a figure that loads a postscript file. You can use the t and b options for selectively putting a figure at the top or bottom of a page. Always include p after t or b so that you won’t risk running out of space if a figure is too big. The h option for inserted figures is not recommended because it’s not used in APA journals.

At any point in the text, before or after the figure, you can automatically refer to it with the ??ommand, for example we can refer to the above figure as Figure 1, where the tilde keeps the word “Figure” (which is capitalized according to APA instructions) together with the reference number. Use names that are easy to remember instead of “fig1” !

Always put the “label command after (and outside) “caption, otherwise you’ll run into trouble with the man mode not finding any labels to refer to.

Figure 1. This is the caption of the figure.

Figure 2. This is the caption of the figure.

Bitmaps

In case of a bitmap eps (as opposed to a graph or line drawing), you probably don’t want it scaled in man mode, for best reproduction, so use “fitbitmap instead:

Table 1 Some numbers that could be experimental data.

Factor 2

Factor 1  Condition A  Condition B
First  586 (231)  649 (255)
2.2  7.5
Second 590 (195) 623 (231)
2.8  2.5

Tables

What about a table? Well, mostly no change from what you are used to, except that you can use “thickline for the top separator to match the APA-journal look. Here’s an example:

In case of a wide table you would use table* instead of table. Here’s an example:

You can get as fancy as you want with tables, just remember not to use any vertical lines, and to group columns with “cline when they are described by a common header (using “multicolumn).

Use figure* instead of figure for figures that are too wide to fit comfortably in a single column.

Section Header

More text can go here.^{1}

1 And this is a footnote, used in just the regular LaTeX way

ODD-NUMBERED PAGE HEADER 3

Table 2 Some more numbers that could be experimental data and are the same as the previous ones, only more of them.

Factor 2

Factor 1  Condition A  Condition B  Condition C  Condition D  Condition E
First  586 (231)  649 (255)  586 (231)  649 (255)  586 (231)
2.2  7.5  2.2  7.5  2.2
Second  590 (195)  623 (231)  590 (195)  623 (231)  590 (195)
2.8  2.5  2.8  2.5  2.8